# Range, not Independence, Drives Modularity in Biological Inspired Representations

### v2

@@ B.2 Full Theoretical Treatment
Let ${\bm{s}}\in\mathbb{R}^{d_{s}}$, ${\bm{z}}\in\mathbb{R}^{d_{z}}$, ${\bm{W}}_{\mathrm{in}}\in\mathbb{R}^{d_{z}\times d_{s}}$, ${\bm{b}}_{\mathrm{in}}\in\mathbb{R}^{d_{z}}$, ${\bm{W}}_{\mathrm{out}}\in\mathbb{R}^{d_{s}\times d_{z}}$, and ${\bm{b}}_{\mathrm{out}}\in\mathbb{R}^{d_{s}}$, with $d_{z}>>d_{s}$. Consider the constrained optimization problem
@@
To show this statement we begin by showing a lemma: that, for a fixed encoding size, the weight loss is minimised by orthogonalising the encoding of each source, the multidimensional analogue of Section B.1.3. Writing the representation as an affine function of the de-meaned sources:
@@
The activity loss difference between these two representations is:
$\mathscr{L}_{G}^{\text{Modular}}-\mathscr{L}_{G}^{\text{Mixed}}=\sum_{j}u_{nj}^{2}(\min_{i}\bar{s}_{j}^{[i]})^{2}-(\min_{i}[\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}])^{2}-\sum_{j=1,k\neq j}^{d_{s}}u_{nj}u_{nk}\langle\bar{s}_{j}^{[i]}\bar{s}_{k}^{[i]}\rangle_{i}$ (48)
@@
So, if these inequalities are satisfied, both the weight and activity loss are minimised at fixed encoding size by modularising, and this holds for all encoding sizes, therefore the optimal solution must be modular. Hence these conditions are sufficient for modular solutions to be optimal.
@@
For intuition on how corner-cutting relates to the above inequalities see the pedagogical derivation in section Section B.1.
@@ B.3 Range-Independent Variables
We will now prove this corollary, that was pedagogically discussed in Section B.1:
@@
In the same setting as Theorem 2.1 the optimal representation modularises if all sources are pairwise extreme-point independent, i.e. if

### v4

@@ B.1 Full Theoretical Treatment
B.1.1 Weight Losses are Minimised by Modularising
First, we show a lemma: that, for a fixed encoding size, the weight loss is minimised by orthogonalising the encoding of each source. Writing the representation as an affine function of the de-meaned sources:
@@ B.1.2 Activity Loss
The activity loss difference between the modular and mixed representations is:
$\sum_{j}u_{nj}^{2}(\min_{i}\bar{s}_{j}^{[i]})^{2}-(\min_{i}[\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}])^{2}-\sum_{j=1,k\neq j}^{d_{s}}u_{nj}u_{nk}\langle\bar{s}_{j}^{[i]}\bar{s}_{k}^{[i]}\rangle_{i}$ (27)
@@
B.1.3 Combination of Weight and Activity
So, if these inequalities are satisfied you can always decrease the loss by modularising a mixed neuron. Therefore, both the weight and activity loss are minimised at fixed encoding size by modularising, and this holds for all encoding sizes, therefore the optimal solution must be modular. Hence these conditions are sufficient for modular solutions to be optimal.
@@ B.1.3 Combination of Weight and Activity
B.2 Equivalence of Convex Hull Formulation
In this section we prove Theorem 2.2, an equivalent formulation of the infinitely many inequalities in Theorem B.1 in terms of the convex hull, restated below:
Theorem B.3.
In the setting as Theorem B.1, define the following matrix:
${\bm{F}}=\begin{bmatrix}(\min_{i}s_{1}^{[i]})^{2}&-\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&-\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{2}^{[i]}\rangle_{i}&(\min_{i}s_{2}^{[i]})^{2}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&\ldots\\ -\langle s_{1}^{[i]}s_{3}^{[i]}\rangle_{i}&-\langle s_{2}^{[i]}s_{3}^{[i]}\rangle_{i}&(\min_{i}s_{3}^{[i]})^{2}&\ldots\\ \vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (29)
Define the set $E=\{{\bm{y}}:{\bm{y}}^{T}{\bm{F}}^{-1}{\bm{y}}=1\}$. Then an equivalent statement of the modularisation inequalities Equation 16, is the representation modularises iff $E$ lies within the convex hull of the datapoints.
Notice that the matrix ${\bm{F}}$ is not guaranteed to be positive definite, meaning $E$ is not necessarily an ellipse. In these cases no convex hull can contain the unbounded quadric it forms, and the optimal solution is mixed regardless of the particular distribution of the datapoints.
B.2.1 Rewrite Inequality Constraints
If this equation holds for one vector, ${\bm{w}}$, it also holds for all positively scaled versions of the vector, ${\bm{v}}=\lambda{\bm{w}}$ for $\lambda>0$, so we can equivalently state that this condition must only hold for unit vectors.
Construct the matrix ${\bm{F}}$ as above Equation 29, then this condition can be rewritten:
$(\min_{i}{\bm{w}}^{T}{\bm{s}}^{[i]})^{2}\geq{\bm{w}}^{T}{\bm{F}}{\bm{w}}\quad\forall{\bm{w}}\in\mathbb{R}^{n},|{\bm{w}}|=1$ (31)
$(\max_{i}{\bm{w}}^{T}{\bm{s}}^{[i]})^{2}\geq{\bm{w}}^{T}{\bm{F}}{\bm{w}}\quad\forall{\bm{w}}\in\mathbb{R}^{n},|{\bm{w}}|=1$ (32)
B.2.2 Sufficiency of Condition
${\bm{y}}_{w}=\operatorname*{arg\,max}_{{\bm{y}}\in E}{\bm{w}}^{T}{\bm{y}}$ (33)
We can find ${\bm{y}}_{w}$ by lagrange optimisation under the constraint that ${\bm{y}}^{T}{\bm{F}}^{-1}{\bm{y}}=1$:
${\bm{y}}_{w}=\frac{{\bm{F}}{\bm{w}}}{\sqrt{{\bm{w}}^{T}{\bm{F}}{\bm{w}}}}$ (34)
Now:
${\bm{w}}^{T}{\bm{y}}=\sqrt{{\bm{w}}^{T}{\bm{F}}{\bm{w}}}=\sum_{i}\lambda_{i}{\bm{w}}^{T}{\bm{s}}^{[i]}\leq\max_{i}{\bm{w}}^{T}{\bm{s}}^{[i]}\sum_{i}\lambda_{i}=\max_{i}{\bm{w}}^{T}{\bm{s}}^{[i]}$ (36)
Therefore, we get our our main result:
$(\max_{i}{\bm{w}}^{T}{\bm{s}}^{[i]})^{2}\geq{\bm{w}}^{T}{\bm{F}}{\bm{w}}$ (37)
B.2.3 Necessity of Condition
First, get all the datapoints that for some unit vector ${\bm{w}}$ maximise ${\bm{w}}^{T}{\bm{s}}^{[i]}$ over the dataset. Construct the convex hull of these points, the polyhedron $P$. You can equivalently characterise any polyhedra by a set of linear inequalities: $P=\{{\bm{x}}\in\mathbb{R}:{\bm{B}}{\bm{x}}\leq{\bm{b}}\}$. Denote with ${\bm{B}}_{j}$ the jth row of ${\bm{B}}$. We can always choose to write the inequalities in such a way that the rows of ${\bm{B}}$ are unit length, by appropriately rescaling each $b_{j}$. Assume we have done this.
Assume the convex hull does not contain the entire ellipse. Therefore there is at least one point on the ellipse that breaks one of the inequalities defining the polyhedron. These points are charactersied by being both on the ellipse, ${\bm{y}}^{T}{\bm{F}}^{-1}{\bm{y}}=1$, but at least one of the inequalities defining the polyhedron is broken: ${\bm{B}}_{j}^{T}{\bm{y}}>b_{j}$. Let’s call the set of points on the ellipse for which this inequality is broken $\mathscr{B}_{j}=\{{\bm{y}}\in\mathbb{R}^{n}:{\bm{y}}^{T}{\bm{F}}^{-1}{\bm{y}}=1\quad\text{and}\quad{\bm{B}}_{j}^{T}{\bm{y}}>b_{j}\}$.
Now, choose the element of $\mathscr{B}_{j}$ that maximises ${\bm{B}}_{j}^{T}{\bm{y}}$, call it ${\bm{y}}^{*}$. This will also be the element of $E$ that maximises ${\bm{B}}_{j}^{T}{\bm{y}}$, which by the previous lagrange optimisation, Equation 34, has the following dot product: ${\bm{B}}_{j}^{T}{\bm{y}}^{*}=\sqrt{{\bm{B}}_{j}^{T}{\bm{F}}{\bm{B}}_{j}}$.
But now we derive a contradiction. This point is outside the polyhedron, ${\bm{B}}_{j}^{T}{\bm{y}}>b_{j}$, whereas all points inside the polyhedron, including all the datapoints, satisfy all inequalities, hence, ${\bm{B}}_{j}^{T}{\bm{s}}^{[i]}\leq b_{j}\forall{\bm{s}}^{[i]}$. But, ${\bm{B}}_{j}^{T}{\bm{y}}^{*}=\sqrt{{\bm{B}}_{j}^{T}{\bm{F}}{\bm{B}}_{j}}>b_{j}$, therefore:
${\bm{B}}_{j}^{T}{\bm{s}}^{[i]}\leq b_{j}<\sqrt{{\bm{B}}_{j}^{T}{\bm{F}}{\bm{B}}_{j}}\quad\forall{\bm{s}}^{[i]}$ (38)
And hence we’ve found a unit vector ${\bm{B}}_{i}$ such that:
$\max_{i}{\bm{B}}_{j}^{T}{\bm{s}}^{[i]}<\sqrt{{\bm{B}}_{j}^{T}{\bm{F}}{\bm{B}}_{j}}\quad\rightarrow\quad(\max_{i}{\bm{B}}_{j}^{T}{\bm{s}}^{[i]})^{2}<{\bm{B}}_{j}^{T}{\bm{F}}{\bm{B}}_{j}$ (39)
@@ B.3 Range-Independent Variables
We will now prove the corollary on range dependence, corollary 2.3, recapped here:
@@
In the same setting as Theorem B.1 the optimal representation modularises if all sources are pairwise extreme-point independent, i.e. if